OpenStudy (anonymous):

Is the statement true or false? n!(n+1)=(n+1)!

OpenStudy (anonymous):

could you help me out?

OpenStudy (anonymous):

may you help me out?

OpenStudy (solomonzelman):

\(\huge {\color{blue} {true} }\)

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

why is it true? could you show me how to work it out?

OpenStudy (jdoe0001):

\(\bf 5!\implies 5\cdot 4\cdot 3\cdot 2\cdot 1\\ \quad \\ (5+1)!\implies 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1\implies 6\cdot 5!\implies 5!\cdot (5+1)\)

OpenStudy (solomonzelman):